2 2 ut Given this information, which of the following is most likely to occur if the actual unemployment in suppose the Phillips curve is represented by the following equation T, any period is equal to 10%? π1-1 O the rate of inflation will be constant O the rate of inflation will tend to increase O the rate of inflation will tend to decrease None of the options listed
Answer #1

Answer :Option A is correct. As when actual unemployment rate equal to 10%

The rate of inflation will be constant as it shows that there is no change in the inflation level.As Difference between two inflation level is zero.

Difference between two period of inflation π​​​​​​ t - π t -1 =20-2(10)=0